# Paper/main_online_appendix

### v3

@@ 6 Relationship to Existing Models
Relationship to Linear Differentiation. He and Natenzon, 2023b propose a Linear Differentiation Model (LDM) defined on the same multiattribute domain as our $L_{1}$-complexity model, where binary choice probabilities are given by
@@
for linear utility $U(x)=\sum_{k}\beta_{k}x_{k}$, an $n\times n$ positive definite matrix $\Sigma$, and a continuous, strictly increasing $G$. Note that this model is identical to the $L_{1}$-complexity model save for the choice of distance metric in the denominator: in the LDM, dissimilarity is measured by the generalized Euclidean distance in contrast to the $L_{1}$ distance. Below, we outline two key formal dimensions along which the models differ.
@@
Recall that in our model, the DM only receives information on ordinal value comparisons. In Bayesian Probit, the DM learns about the cardinal value differences between choice options, which rules out certain intuitive choice patterns. For instance, consider the the following choice options in the multiattribute domain:
$x$ $=(2,0)$
$y$ $=(0,1)$
$z$ $=(0,0).$
[…]
@@
Importantly, this feature of Bayesian Probit means it cannot capture key choice patterns in several of the applications considered in this paper. For instance, in the application to lottery valuation, we model settings where the DM finds a given lottery trivial to compare against certain options in the price list and difficult to compare to others — for instance, the DM is certain that the lottery that pays \$10 with 80% chance is worse than \$10 for sure, yet finds the same lottery difficult to compare to \$5 for sure — and yet perfectly understands the ranking within the price list, i.e. that \$10 is better than \$5. Bayesian Probit cannot accomodate these choice patterns, whereas our choice model can.
@@ 7 Conclusion
This paper presents a theory of tradeoff-based comparison complexity in multiattribute, lottery, and intertemporal choice, in which comparisons are easier when options are more similar along their features, holding fixed their value difference fixed; and easiest when two options have a dominance relationship. We show how our theory can organize a range of documented biases and instabilities in choice that operate across different choice domains, and generates novel predictions on how these patterns can be eliminated or even reversed by manipulating the nature of tradeoffs. We test the validity of our theory of comparison complexity using large-scale binary choice data, and run additional experiments to test the novel predictions of our model for choice and valuation. Finally, we apply our model to develop insights on strategic obfuscation by firms in a pricing game.
APPENDIX
@@ Appendix A Appendix: Tables and Figures
Dependent Variable: Error Rate Dependent Variable: Inconsistency Rate Dependent Variable: CU
@@
Global Value Difference $0.00$ $-0.01$ $-0.00$
@@
OLS Estimates. Standard errors (in parentheses) are robust. “$L_{1}$ Ratio" and “Global Value Difference" are the $L_{1}$ ratio and the monetary value difference for each choice problem. ${}^{***}p<0.001$; ${}^{**}p<0.01$; ${}^{*}p<0.05$.
@@
Dependent Variable: Error Rate Dependent Variable: Inconsistency Rate Dependent Variable: CU
@@
Global Value Difference $-0.00$ $-0.01$ $-0.00^{*}$
@@
OLS Estimates. Standard errors (in parentheses) are robust. “$L_{1}$ Ratio" and “Global Value Difference" are the $L_{1}$ ratio and the monetary value difference for each choice problem. ${}^{***}p<0.001$; ${}^{**}p<0.01$; ${}^{*}p<0.05$.
@@
“DF”, “BGS”, “Focus”, and “RT” refer to the Distortion-Free, Salience, Focusing, and Relative Thinking models described in Appendix E.1. “$L_{1}$-C, 2P” and “$L_{1}$-C, 3P” refer to the 2 and 3 parameter $L_{1}$-Complexity models described in Appendix E.1. Completeness and Restrictiveness measures are defined in Appendix F. Standard errors for Restrictiveness estimates in parentheses.
@@
Dependent Variable: Error Rate Dependent Variable: Inconsistency Rate Dependent Variable: CU
@@
Global Value Difference $-0.00$ $-0.01^{***}$ $-0.00$
@@
Dependent Variable: Binary Error (Indiv. $\hat{\delta}$)
@@
Indiv. Value Difference $-0.01^{***}$ $-0.01^{***}$
@@
Dependent Variable: Error Rate Dependent Variable: Inconsistency Rate Dependent Variable: CU
@@
Global Value Difference $-0.01^{***}$ $0.00^{***}$ $0.01^{***}$
@@
Dependent Variable: Binary Error (Indiv. $\hat{\alpha}$)
@@
Indiv. Value Difference $-0.00$ $0.00$
@@
   $\eta$ 0.22 0.24 0.33
$R^{2}$ 0.54 0.55 0.59 0.65 0.72
Completeness 0.64 0.66 0.71 0.81 0.9
Restrictiveness 0.564 0.564 0.532 0.6 0.596
(0.000) (0.000) (0.000) (0.000) (0.000)
“EU", “RDEU", and “CPT" refer to the Expected Utility, Reference-Dependent Expected Utility, and Cumulative Prospect Theory models described in Appendix E.3. “EV CDF-C" and “EU CDF-C" refer to the risk-neutral and expected utility CDF complexity models described in Appendix E.3. Completeness and Restrictiveness measures are defined in Appendix F. Standard errors for Restrictiveness estimates in parentheses.
@@ Appendix B Appendix: Additional Theoretical Results
Proofs of all results stated in this Appendix are compiled in Online Appendix F.
@@ B.1.2 Multiattribute Choice: Nonlinear Preferences
We consider an extended multiattribute domain where each option in $X\equiv X_{1}\times X_{2}\times...\times X_{n}$ is defined on $n$ attribute dimensions, where each $X_{i}$ is a connected and separable topological space. Preferences are additively separable in each attribute, where the value of each $x$ is given by $U(x)=\sum_{k}u_{k}(x_{k})$ for $u_{k}$ continuous. Say that $u_{k}$ is non-trivial if there exist $x_{k},x^{\prime}_{k}\in X_{k}$ such that $u_{k}(x_{k})\neq u_{k}(x^{\prime}_{k})$. We propose that the ease of comparison in this domain is governed by the following representation:
@@
Note that this representation for $\tau$ satisfies the same principles of similarity, dominance, and simplification as the linear representation introduced in Section 2.1, and the corresponding representation for $\rho(x,y)$ satisfies axioms M1 and M3–M5. We provide an axiomatic characterization for this representation, in which Linearity is relaxed and replaced with two axioms.
@@
Separability is the stochastic analog of the familiar coordinate independence axiom in deterministic choice, which says that $x_{E}z\succeq y_{E}z\implies x_{E}z^{\prime}\succeq y_{E}z^{\prime}$ for all $E\subseteq I$, $x,y,z,z^{\prime}\in X$. The interpretation of Tradeoff Congruence is as follows: consider the attribute-wise tradeoffs involved in comparing $z$ to $y$ and $x$ to $y$, where $x$ is in fact better than $y$, and $y$ is better than $z$. If replacing $y$ with $x$ in the first comparison and replacing $y$ with $z$ in the second only increases the magnitude of these tradeoffs – i.e. if $(x,y)$ and $(y,z)$ are congruent – then $(x,z)$ cannot be an easier comparison than both of the intermediate comparisons $(x,y)$ and $(y,z)$. Intuitively, both of these replacements only increase the size of the tradeoffs the DM must contend with, and so as revealed by choice probabilities, the DM cannot find the comparison $(x,z)$ easier than both $(x,y)$ and $(y,z)$.
@@ B.1.3 Mulitrattribute Choice: Specifying Attributes
Suppose $\rho$ is represented by $(G,P,(u_{E})_{E\in P})$ and $(\tilde{G},\tilde{P},(\tilde{u}_{E})_{E\in\tilde{P}})$. Then $G=\tilde{G}$ and $P=\tilde{P}$, and there exists $C>0$, and $b_{E}\in\mathbb{R}$ such that for each $E\in P$, $\tilde{u}_{E}=Cu_{E}+b_{E}$.
@@ B.1.5 Intertemporal Choice
[…]
@@ B.3 Tiebreaking in Multinomial Choice
$\mathcal{C}(x,s)\equiv\begin{cases}1/\mathcal{N}(x,s)&\mathbb{E}[v_{x}|s]\geq\mathbb{E}[v_{y}|s]\,\forall\,y\in A\\ 0&\text{otherwise}\end{cases}$
@@
$\mathcal{C}^{i}(x,s)\equiv\begin{cases}1/\mathcal{N}^{i}(x,s)&\mathbb{E}[v_{x}|s]\geq\mathbb{E}[v_{y}|s]\,\forall\,y\in A^{i}\\ 0&\text{otherwise}\end{cases}$
@@
$\rho((x^{1},...,x^{n}),(A^{1},...,A^{n})|C)=\mathbb{E}\left[\prod_{i=1}^{n}\mathcal{C}^{i}(x^{i},s)\,\bigg{|}\,v\right].$
@@ B.4 Identification in Multinomial Choice
Call $\rho:X\times\mathcal{A}\times\mathcal{M}\to[0,1]$ a multinomial choice rule if $\sum_{x\in A}\rho(x,A|C)=1$ for all $(A,C)\in\mathcal{A}\times\mathcal{M}$. Our multinomial choice model is parameterized by the prior distribution $Q$, the value function $v:X\to\mathbb{R}$, and the signal precisisons $\tau:\mathcal{D}\to\mathbb{R}^{+}$, where we make the additional assumption that $\tau(x,y)=0$ if $v(x)=v(y)$. The following result states that $v$ is ordinally identified and $\tau$ is exactly identified.
@@ B.5 Decoy Effects
Comparison to other context-dependent models. Though each of the choice patterns above can be explained by familiar models, our model is distinct in simultaneously explaining all three. The salience (Bordalo et al.,, 2013) and focusing models (Kőszegi and Szeidl,, 2013) cannot rationalize the decoy effects in Cases 1 and 2. The relative thinking model (Bushong et al.,, 2021), in which the DM weighs a given change along an attribute by less when there is a larger range of values along that attribute, can rationalize the decoy effect in Case 1 as a result of option $z$ extending the range of attribute 2 more than attribute 1, but not Case 2, where $z^{\prime}$ has no effect on attribute ranges. The pairwise normalization model (Landry and Webb,, 2021) predicts that $z$ increases the relative of $y$ relative to $x$ whenever $z_{1}/z_{2}$ is closer to $y_{1}/y_{2}$ than it is to $x_{1}/x_{2}$, and so can rationalize the decoy effects in both Cases 1 and 2, but also delivers the counterfactual prediction that the addition of a mutually dominated option $z^{\prime\prime}$ will also distort choice in favor of $x$. Furthermore, all of these models are formulated in multi-attribute choice, which means they cannot easily explain documented decoy effects in lottery choice (Soltani et al.,, 2012) or other domains. Our choice framework straightforwardly applies to lottery and intertemporal choice.
We also make an important conceptual distinction from these existing models. In our model, biased choice does not arise from a behavioral bias, but instead as a rational response to imperfect comparability. We only expect decoy options to distort choice between $x$ and $y$ when their binary comparison is challenging. This is consistent with the fact that the attraction effect is muted when consumers face familiar choice contexts or have clear prior preferences (Huber et al.,, 2014) – an empirical phenomenon that is not well-explained by existing models.

### v4

@@ 5 Relationship to Existing Models
Relationship to Linear Differentiation. He and Natenzon, (2023) propose a Linear Differentiation Model (LDM) in multiattribute choice, where choice probabilities are given by
@@
for linear utility $U(x)=\sum_{k}\beta_{k}x_{k}$, an $n\times n$ positive definite matrix $\Sigma$, and a continuous, strictly increasing $G$. Note that this model is identical to the $L_{1}$-complexity model save for the choice of distance metric in the denominator: in the LDM, dissimilarity is measured by the generalized Euclidean distance, in contrast to the $L_{1}$ distance. Below, we outline two key formal dimensions along which the models differ.
@@
Recall that in our model, the DM only receives information on ordinal value comparisons. In Bayesian Probit, the DM learns about the cardinal value differences between choice options, which rules out certain intuitive choice patterns. For instance, consider the the following choice options in the multiattribute domain: $x=(2,0),y=(0,1),z=(0,0)$.
Since $z$ is dominated by both $x$ and $y$, we might expect that $\rho(x,z)=\rho(y,z)=1$ and also $\rho(x,y)<1$; that is, the DM does not err in the presence of dominance but finds tradeoffs across attributes difficult. [^33] The $L_{1}$-complexity model generates the choice probabilities for $\beta_{1},\beta_{2}$¿0, and G(1)=G(1)=1G(-1)=G(1)=1. Bayesian Probit cannot rationalize this choice data; $\rho(x,z)=\rho(y,z)=1$ implies that $\operatorname{\text{Cor}}(\epsilon_{x},\epsilon_{z})=\operatorname{\text{Cor}}(\epsilon_{y},\epsilon_{z})=1$, which implies $\operatorname{\text{Cor}}(\epsilon_{x},\epsilon_{y})=1$; as a result, we have $\rho(x,y)=1$.343434 More generally, when $\rho(x,z),\rho(y,z)$ are close but not necessary equal to 1, the Bayesian Probit model places lower bounds on the choice probability $\rho(x,y)$. As Appendix B.10 discusses, these bounds imply that there are binary choice rules rationalizable by $L_{1}$, CDF, and CPF complexity but not by Bayesian Probit.
@@
Importantly, this feature of Bayesian Probit means it cannot capture key choice patterns in several of the applications considered in this paper. For instance, in the application to lottery valuation, we model settings where the DM finds a given lottery trivial to compare against certain options in the price list and difficult to compare to others — for instance, the DM is certain that the lottery that pays \$10 with 80% chance is worse than \$10 for sure and finds the same lottery difficult to compare to \$5 for sure — yet perfectly understands the ranking within the price list, i.e. that \$10 is better than \$5. Bayesian Probit cannot accommodate these choice patterns, whereas our choice model can.
@@ 6 Conclusion
This paper presents a theory of tradeoff-based comparison complexity in multiattribute, lottery, and intertemporal choice. We show how our theory can be deployed to organize a range of documented biases and choice instabilities across different domains of behavioral economics, and generates novel predictions on how these patterns can be eliminated or reversed by manipulating the nature of tradeoffs. We demonstrate the predictive power of our complexity measures using large-scale binary choice data, and experimentally test the novel predictions of our theory. Here, we close by discussing limitations of the paper and potential avenues for future work.
Incorporating systematic biases. This paper shows that a parsimonious theory of noise in binary comparisons can rationalize systematic distortions that resemble commonly modeled biases. Throughout, we abstract away from potential biases in the DM’s objective function $v_{x}$. However, even biased decision-makers presumably face tradeoffs, and so there may be value in integrating models of systematic biases — that is, models that augment $v_{x}$ — with our model of tradeoff-driven complexity, and understanding their behavioral implications. Some models can be easily integrated with our approach: for instance, many non-standard multiattribute choice models simply reflect distorted attribute weights, which can directly be incorporated into the $L_{1}$ complexity measure. However, incorporating other distortions may require more involved adaptations of our measure.
Developing implications for measurement. This paper develops a theory that can help explain apparent inconsistencies across preference elicitation methods, such as the gap between certainty equivalents and binary choice found in lottery preference reversals. However, we do not view our theory nor experimental results as an indication that any one elicitation method should be a priori favored over others — the suitability of a given measure ultimately depends on the goals of the researcher and how response data is interpreted in service of those goals. [^35] For instance, even if valuations are subject to pull-to-center effects as in our model, comparing the valuations of two options could still be informative about option the DM prefers, so long as pull-to-center effects impact valuations in the same way for both options. Likewise, inference of preference intensities on the basis of binary choice rates may be counfounded by the presence of heteroskedastic noise; see McGranaghan et al., (2024) for an illustration of this idea. Instead, we view a theory like ours, which models how noise may manifest differently across elicitation methods, as a tool to help facilitate the interpretation of existing elicitation methods and motivate the development of new measures. We view both aims as promising directions for future work.
Extending to additional domains. Although we formulated our theory in the domains of multiattribute, lottery, and intertemporal choice, the difficulty of tradeoffs is likely a more general feature of decision-making. One straightforward extension is to adapt our $L_{1}$ complexity measure to choice under uncertainty by reinterpreting the attributes as states of the world. Another potential extension is to apply the model to derive solution concepts in normal-form games that account for tradeoff-induced noise.
Heterogeneity in tradeoff difficulty. Our theory implies that agents with identical preferences will agree on the relative difficulty of comparisons, since the value dissimilarity ratio is pinned down by the DM’s preference parameters. However, these agents could differ in the absolute difficulty they face in a given comparison — there could be heterogeneity in the $G$ transformation that maps the value-dissimilarity ratio into binary choice rates. Intuitively, decision-makers may differ not only along their preferences, but also in how they respond to complexity, for instance due to differences in cognitive ability or attentional costs. Measuring this latter form of heterogeneity could allow researchers to better understand correlations (or lack thereof) across behavioral regularities and preference measures.
@@ Appendix A Appendix: Tables and Figures
Dependent Variable:
Error Rate
Dependent Variable:
Inconsistency Rate
Dependent Variable:
CU
@@
Global Value Difference $0.00$ 0.01-0.01 0.00-0.00
@@
OLS Estimates. Standard errors (in parentheses) are robust. $L_{1}$ Ratio" and “Global Value Difference" are the $L_{1}$ ratio and the monetary value difference for each choice problem. $^{***}p<0.001$; $^{**}p<0.01$; $^{*}p<0.05$.
@@
Dependent Variable:
Error Rate
Dependent Variable:
Inconsistency Rate
Dependent Variable:
CU
@@
Global Value Difference 0.00-0.00 0.01-0.01 $0.00^{*}$
@@
OLS Estimates. Standard errors (in parentheses) are robust. $L_{1}$ Ratio" and “Global Value Difference" are the $L_{1}$ ratio and the monetary value difference for each choice problem. $^{***}p<0.001$; $^{**}p<0.01$; $^{*}p<0.05$.
@@
“DF”, “BGS”, “Focus”, and “RT” refer to the Distortion-Free, Salience, Focusing, and Relative Thinking models described in Appendix E.1. $L_{1}$-C, 2P” and $L_{1}$-C, 3P” refer to the 2 and 3 parameter $L_{1}$-Complexity models described in Appendix E.1. Completeness and Restrictiveness measures are defined in Appendix F. Standard errors for Restrictiveness estimates in parentheses.
@@
Dependent Variable:
Error Rate
Dependent Variable:
Inconsistency Rate
Dependent Variable:
CU
@@
Global Value Difference 0.00-0.00 $1^{***} 0.00-0$.00
@@
Dependent Variable:
Binary Error (Indiv. $\hat{\delta}$)
@@
Indiv. Value Difference $02^{***} 0.0$$01^{***}$
@@
Dependent Variable:
Error Rate
Dependent Variable:
Inconsistency Rate
Dependent Variable:
CU
@@
Global Value Difference $1^{***} 0.000$$^{***} 0.0$$01^{***}$
@@
Dependent Variable:
Binary Error (Indiv. $\hat{\alpha}$)
@@
Indiv. Value Difference 0.00-0.00 $0.00$
@@
   $\eta$ 0.22 0.24 0.24 0.33
$R^{2}$ 0.54 0.56 0.55 0.59 0.65 0.72
Completeness 0.64 0.67 0.66 0.71 0.81 0.9
Restrictiveness 0.563 0.559 0.563 0.533 0.6 0.596
(0.000) (0.000) (0.000) (0.000) (0.000) (0.000)
“EU”, “ST”, “RDEU”, and “CPT” refer to the Expected Utility, Simplicity Theory, Reference-Dependent Expected Utility, and Cumulative Prospect Theory models described in Appendix E.3. “EV CDF-C” and “EU CDF-C” refer to the risk-neutral and expected utility CDF complexity models described in Appendix E.3. Completeness and Restrictiveness measures are defined in Appendix F. Standard errors for Restrictiveness estimates in parentheses.
@@ Appendix B Appendix: Additional Theoretical Results
Proofs of all results stated in this Appendix are compiled in Supplemental Appendix F.
@@ B.1.2 Multiattribute Choice: Nonlinear Preferences
We consider a more general multiattribute domain. Each option in $X\equiv X_{1}\times X_{2}\times...\times X_{n}$ is defined on $n$ attributes, where each $X_{i}$ is a connected and separable topological space. Preferences are additively separable in each attribute, where the value of each $x\in X$ is given by $U(x)=\sum_{k}u_{k}(x_{k})$. Say that $u_{k}$ is non-trivial if there exist $x_{k},x^{\prime}_{k}\in X_{k}$ such that $u_{k}(x_{k})\neq u_{k}(x^{\prime}_{k})$. We propose that the ease of comparison in this domain is governed by the following representation:
@@
Note that this representation mirrors the linear $L_{1}$-complexity representation discussed in the main text, except each utility-weighted attribute value $\beta_{k}x_{k}$ is replaced by its potentially non-linear counterpart $u_{k}(x_{k})$. We provide an axiomatic characterization for this more general representation, in which Linearity is relaxed and replaced with two axioms.
@@
Separability is the stochastic analog of the familiar coordinate independence axiom in deterministic choice, which says that $x_{E}z\succeq y_{E}z\implies x_{E}z^{\prime}\succeq y_{E}z^{\prime}$ for all $E\subseteq I$, $x,y,z,z^{\prime}\in X$. The interpretation of Tradeoff Congruence is as follows: consider the attribute-wise tradeoffs involved in comparing $z$ to $y$ and $x$ to $y$, where $x$ is in fact better than $y$, and $y$ is better than $z$. The condition says that if replacing $y$ with $x$ in the first comparison and replacing $y$ with $z$ in the second only increases the magnitude of these tradeoffs — i.e. if $(x,y)$ and $(y,z)$ are congruent — then $(x,z)$ cannot be an easier comparison than both of the intermediate comparisons $(x,y)$ and $(y,z)$. Intuitively, both of these replacements only increase the size of the tradeoffs the DM must contend with, and so as revealed by choice probabilities, the DM cannot find the comparison $(x,z)$ easier than both $(x,y)$ and $(y,z)$.
@@ B.1.3 Mulitrattribute Choice: Specifying Attributes
[…]
Supplemental Appendix J provides axiomatic foundations for the generalized $L_{1}$-complexity representation.
@@ B.1.5 Intertemporal Choice
Note that if $d$ is differentiable, $d_{CPF}$ can be rewritten as $d_{CPF}(x,y)=\int_{0}^{\infty}|M_{x}(t)-M_{y}(t)|\cdot(-d^{\prime}(t))\,dt$. In the case where $d(t)=\delta^{t}$, generalized CPF complexity reduces to Definition 4. Let $\geq$ denote the partial order $X$ corresponding to temporal dominance (i.e., $x\geq y$ iff at every time $t\in\mathbb{R}^{+}\cup\{+\infty\}$, $M_{x}(t)\geq M_{Y}(t))$. Consider the following axioms:
@@ B.3 Tiebreaking in Multinomial Choice
$c(x,s)\equiv\begin{cases}1/\mathcal{N}(x,s)&\mathbb{E}[v_{x}|s]\geq\mathbb{E}[v_{y}|s]\,\forall\,y\in A\\ 0&\text{otherwise}\end{cases}$
@@
$c^{i}(x,s)\equiv\begin{cases}1/\mathcal{N}^{i}(x,s)&\mathbb{E}[v_{x}|s]\geq\mathbb{E}[v_{y}|s]\,\forall\,y\in A^{i}\\ 0&\text{otherwise}\end{cases}$
@@
$\rho((x^{1},...,x^{n}),(A^{1},...,A^{n})|C)=\mathbb{E}\left[\prod_{i=1}^{n}c^{i}(x^{i},s)\,\bigg{|}\,v\right].$
@@ B.4 Identification in Multinomial Choice
Call $\rho:X\times\mathcal{A}\times\mathcal{C}\to[0,1]$ a multinomial choice rule if $\sum_{x\in A}\rho(x,A|C)=1$ for all $(A,C)\in\mathcal{A}\times\mathcal{C}$. Our multinomial choice model is parameterized by the prior distribution $Q$, the value function $v:X\to\mathbb{R}$, and the signal precisisons $\tau:\mathcal{D}\to\mathbb{R}^{+}$, where we make the additional assumption that $\tau(x,y)=0$ if $v(x)=v(y)$. The following result states that $v$ is ordinally identified and $\tau$ is exactly identified.
@@ B.5 Decoy Effects
Comparison to other context-dependent models. Though each of the choice patterns above can be explained by familiar models, our model is distinct in simultaneously explaining all three. The salience (Bordalo et al.,, 2013) and focusing models (Kőszegi and Szeidl,, 2013) cannot rationalize the decoy effects in Cases 1 and 2. The relative thinking model (Bushong et al.,, 2021), in which the DM weighs a given change along an attribute by less when there is a larger range of values along that attribute, can rationalize the decoy effect in Case 1 as a result of option $z$ extending the range of attribute 2 more than attribute 1, but not Case 2, where $z^{\prime}$ has no effect on attribute ranges. The pairwise normalization model (Landry and Webb,, 2021) predicts that $z$ increases the relative of $y$ relative to $x$ whenever $z_{1}/z_{2}$ is closer to $y_{1}/y_{2}$ than it is to $x_{1}/x_{2}$, and so can rationalize the decoy effects in both Cases 1 and 2, but also delivers the prediction that the addition of a mutually dominated option $z^{\prime\prime}$ will also distort choice in favor of $x$. Furthermore, all of these models are formulated in multi-attribute choice, which means they cannot easily explain documented asymmetric dominance/decoy effects in lottery (Soltani et al.,, 2012) or intertemporal (Marini and Paglieri,, 2019) choice. Our choice framework straightforwardly applies to lottery and intertemporal choice.
We also make a conceptual distinction from these existing models. In our model, biased choice does not arise from a mechanical bias, but instead as a rational response to imperfect comparability. We only expect decoy options to distort choice between $x$ and $y$ when their binary comparison is challenging. This is consistent with the fact that the attraction effect is muted when consumers face familiar choice contexts or have clear prior preferences (Huber et al.,, 2014) – a phenomenon that is not well-explained by existing models.